The information presented here represents selected data from the December 31, 2016, balance sheets and income statements for the year then ended for three firms.

Not sure I am on the right track below, help?

Required:

Calculate the missing amounts for each firm.

The information presented here represents selected data from the December 31, 2016, balance sheets and income statements for the year then ended for three firms Required Calculate the missing amounts for each firm Firm A Firm B Firm C S 400,000 532,000 321,000 Total assets, 12/31/16 Total liabilities, 12/31/16 Paid-in capital, 12/31/16 Retained earnings, 12/31/16 Net income for 2016 Dividends declared and paid during 2016 Retained earnings, 1/1/16 215,000 71,000 114,000 146,000 84,000 302,000 87,000 27,000 46,000 117,000 64,000 46.000 48,000 75.000
Answer #1

Firm A Firm B Firm Total assets, 12/31/16 Total liabilities, 12/31/16 Paid - in capital, 12/31/16 Retained earnings, 12/31/16

Firm A :1/1/2016 Retained earning + Net income - Dividend = Retained earning 12/31/16

75,000 + Net income - 48,000 = 114,000

Net Income + 27,000 = 114,000

  Net Income = 114,000 - 27,000

Firm A : Net Income = $87,000

Firm B : 1/1/2016 Retained earning + Net income - Dividend = Retained earning 12/31/16

Retained earning + 87,000 - 27,000 = 302,000

Retained earning + 60,000 = 302,000

Retained Earning = 302,000 - 60,000

1/1/2016 Retained earning = $242,000

Firm C: 1/1/2016 Retained earning + Net income - Dividend = Retained earning 12/31/16

46,000 + 117,000 - 64,000 = Retained earning 12/31/16

163,000 - 64,000 = Retained earning 12/31/16

Retained earning 13/31/16 = $99,000

Assets = Liability + paid in capital + Retained earnings 12/31/16

$321,000 = Liability + 46,000 + 99,000

321,000 = Liability + 145,000

Liability = 321,000 - 145,000

Liability = $176,000